OpenStudy (anonymous):

finding intercepts, x²+3x+1=0

hero (hero):

Use \[ac = (\frac{b}{2}+d)(\frac{b}{2}-d) \] to split 3x into two middle middle terms the factor by grouping

OpenStudy (anonymous):

ah, thank you!

hero (hero):

:o)

jimthompson5910 (jim_thompson5910):

x-intercepts are \[\large \left(\frac{-3+\sqrt{5}}{2},0\right),\left(\frac{-3-\sqrt{5}}{2},0\right)\]. See attached image for how I got the x-intercepts. y-intercept can be found by plugging in x=0 and simplifying

hero (hero):

I hate when people use programs to post solutions.

hero (hero):

I have programs I could use to post solutions, but I don't use them. It's classless.

jimthompson5910 (jim_thompson5910):

nothing wrong with it, just another way to show the answer (plus the steps)

hero (hero):

You can post the steps yourself. You don't have to use silly programs.

OpenStudy (anonymous):

easy, easy. its alright, he uses a program, you dont. and thats okay.

jimthompson5910 (jim_thompson5910):

Yes I could, but I've done it a billion times, so I figured a program helping me out once in a while couldn't hurt.

hero (hero):

No, then it becomes someone else's work, and no tyour own.

jimthompson5910 (jim_thompson5910):

it's my program. I wrote it myself

jimthompson5910 (jim_thompson5910):

it's in perl

hero (hero):

you can't prove that.

jimthompson5910 (jim_thompson5910):

well with a job this specific, it kinda proves itself

OpenStudy (anonymous):

why does it matter if he did or didnt? you both are helping people become confident in their academics, and to me, thats all that matters. so spat if you will, but im equally as grateful.

jimthompson5910 (jim_thompson5910):

besides, if it helps the student, then it shouldn't matter, should it?

OpenStudy (anonymous):

^^^^^

jimthompson5910 (jim_thompson5910):

thx emma, glad to be of help. let me know if I need to improve the explanation or translate anything

OpenStudy (anonymous):

it was perfect, thank you both again so much!